# ABPT: Amended Backpropagation through Time with Partially Differentiable Rewards

### v1

@@ Abstract
Using the exact gradients of the rewards to directly optimize policy parameters via backpropagation-through-time (BPTT) enables high training performance for quadrotor tasks. However, designing a fully differentiable reward architecture is often challenging. Partially differentiable rewards will result in biased gradient propagation that degrades training performance. To overcome this limitation, we propose Amended Backpropagation-through-Time (ABPT), a novel approach that mitigates gradient bias while preserving the training efficiency of BPTT. ABPT combines 0-step and N-step returns, effectively reducing the bias by leveraging value gradients from the learned Q-value function. Additionally, it adopts entropy regularization and state initialization mechanisms to encourage exploration during training. We evaluate ABPT on four representative quadrotor flight tasks. Experimental results demonstrate that ABPT converges significantly faster and achieves higher ultimate rewards than existing learning algorithms, particularly in tasks involving partially differentiable rewards.
@@ 1 Introduction
Reinforcement learning (RL) has shown great success in training control policies and decision-making models, such as quadruped locomotion Kohl and Stone (2004), mobile robots Smart and Kaelbling (2002), and even e-games Vinyals et al. (2019); Hafner et al. (2023). Despite its adaptability, RL relies on zero-order gradient (ZOG) approximations Sutton and Barto (2018) which require extensive sampling or replay mechanisms, which often results in slow convergence and sub-optimal training outcomes. As a promising approach, differentiable physics is considered to be able to accelerate training by providing first-order gradient (FOG) Mozer (2013) with respect to policy parameters. Recent research Xu et al. (2022) has made significant progress in overcoming inherent drawbacks including the tendency to be trapped in local minima or training instability resulting from non-smooth transition in contact-rich environments, which has been demonstrated in several RL benchmarks.
Recent studies on robotics demonstrate that directly using first-order gradients for policy learning achieves faster convergence and superior performance compared to traditional RL algorithms in applications such as quadrotors Zhang et al. (2024); Wiedemann et al. (2023), robotic arms Lv et al. (2023), and quadruped dogs Song et al. (2024). Among these, quadrotors stand out as an ideal platform for policy learning via differentiable physics. This is largely due to their simple and smooth aerodynamics in the flight.
@@
However, using first-order gradients for training requires not only the state transition function but also the reward function to be differentiable. While this condition is manageable for simple tasks, it becomes highly challenging for complex task objectives. Reward functions in such scenarios often involve non-differentiable components, such as conditional constants or binary rewards (e.g., awarding points upon task success), which violate differentiability requirements. These non-differentiable elements disrupt the computation graph during backpropagation-through-time (BPTT), leading to biased first-order gradients. Notably, while first-order gradients remain unbiased under purely non-smooth dynamics Zhang et al. (2023); Suh et al. (2022), they become biased when using non-differentiable rewards—a phenomenon we term Biased Gradient. This bias misguides training, causing optimization to stall in local minima and deviate from the intended direction of improvement.
We propose an on-policy actor-critic approach - Amended Backpropagation-through-Time (ABPT) - to mitigate the bias gradient issue introduced by the non-differentiable rewards while keeping high policy learning performance in terms of training speed and converged rewards. Our approach combines 0-step returns with N-step returns Sutton and Barto (2018), leveraging value gradients generated by the 0-step returns to balance first-order gradient accuracy and exploitation. Additionally, ABPT incorporates entropy to counteract the excessive aggression of first-order gradient. It also employs a replay buffer to store state experiences, initializing episodes with these states to enhance sampling efficiency.
@@
We evaluate our method on four representative quadrotor tasks, comparing it against classic policy gradient and first-order gradient methods. These tasks are designed to progressively increase the reward non-differentiability, testing the adaptability of each approach. Experimental results demonstrate that ABPT achieves the fastest convergence and highest final rewards across all baselines. This superiority is attributed to ABPT’s ability to compensate for biased gradients and enhance exploration via entropy regularization and state replay. Furthermore, ABPT exhibits robustness across varying learning rates and reward structures.
Our technical contributions are summarized as follows:
@@
  • We introduce four representative quadrotor tasks as novel benchmark tests for evaluating different learning approaches.
@@ 2 Related Work
Though first-order gradients enables faster and more accurate gradient computation, accelerating policy training via BPTT Mozer (2013), they suffer from gradient explosion/vanishing or instability caused by non-smooth dynamics. Many attempts try to address these issues and strengthen robustness. PODS Mora et al. (2021) leverages both first- and second-order gradients with respect to cumulative rewards. SHAC Xu et al. (2022) employs an actor-critic framework, truncates the learning window to avoid vanishing/exploding gradients, and smooths the gradient updates. AOBG Suh et al. (2022) combines ZOG (policy gradient) with FOG, using an adaptive ratio based on gradient variance in the minibatch to avoid the high variance typical of pure FOG in discontinuous dynamics. AGPO Gao et al. (2024) replaces ZOG in mixture with critic predictions, as Q-values offer lower empirical variance during policy rollouts. While both AGPO and AOBG converge to asymptotic rewards in significantly fewer timesteps, the mixture ratio requires excessive computational resources, leading to longer wall-time.
@@
Among these algorithms, SHAC Xu et al. (2022) achieves training efficiency comparable to the naive first-order approach (BPTT). However, SHAC primarily focuses on challenges related to dynamics and overlooks the gradient bias introduced by partially differentiable rewards, which are common in real-world quadrotor tasks. These tasks often involve differentiable dynamics but rely on discrete or sparse reward signals, such as waypoint reaching events or minimum-time completions. Our approach directly tackles this limitation by introducing a 0-step return to address gradient bias from non-differentiable rewards and incorporating entropy to encourage exploration in challenging scenarios. The results highlight our method’s superior performance in quadrotor applications, achieving the fastest convergence and the highest final rewards among all the evaluated methods.
3 Methodology
3.1 Preliminaries
@@
Gao et al. (2024) named this gradient estimator as Q gradient (QG). Compared to ZOG, the performance of value approximation is extremely crucial to actor training. This is because QG directly relies on the back-propagation of the action-value function, while ZOG firstly estimates advantages among values according to current policy that enables actor training more robust to the critic.
3.2 First-Order Gradient & Gradient Bias
@@
where $N$ represents the horizon length and $i$ denotes the $i$-th trajectory within the minibatch. To consider infinite return whiling avoiding gradient explosion, an approximated N-step return Sutton and Barto (2018) has been introduced in Xu et al. (2022):
$\small\nabla_{\theta}\mathcal{J}_{\theta}=\left(\sum_{k=0}^{N-1}\gamma^{k}\frac{\partial R(s_{t+k})}{\partial\theta}\right)+\gamma^{N}\nabla_{\theta}V_{\phi}(s_{t+N}).$ (7)
Here, $V_{\phi}$ is the state-value function reparameterized by $\phi$. As shown in Xu et al. (2022), using this approximated N-step return can introduce smooth landscape for optimization and mitigate the gradient explosion issues. However, it cannot address non-differentiable rewards as we will discuss later.
@@
$\small\mathcal{J}_{\theta}=\left(\sum_{k=0}^{N-1}\gamma^{t}\Big{(}R_{dist}(s_{t+k})+R_{succ}(s_{t+k})\Big{)}\right)+\gamma^{N}V_{\phi}(s_{t+N}),$ (8)
3.3 The Proposed Method
@@
$\small\mathcal{J}_{\theta}^{N}=\underbrace{\left(\sum_{k=0}^{N-1}\gamma^{k}{R}(s_{t+k}^{i})\right)}_{\mathcal{G}^{t:t+N}_{\theta}}+(1-d)\gamma^{N}\underbrace{V_{\phi}(s_{t+N}^{i})}_{\mathcal{V}_{\theta|\phi}^{t+N+1}}$ (10)
and
$\mathcal{J}_{\theta}^{0}=\underbrace{V_{\phi}(s_{t}^{i})}_{\mathcal{V}_{\theta|\phi}^{t}}$ (11)
@@
respectively. Here, $d$ is a boolean variable indicating whether the current episode has ended, and $i$ denotes the trajectory index. Given the trajectory is perform by $\pi_{\theta}$, each element is differentiable to parameter $\theta$. $\mathcal{G}_{\theta}^{t:t+N}$ represents the accumulated reward within the horizon and $\mathcal{V}_{\theta|\phi}^{t+N+1}$ is the value obtained by fixed critic.
$\small\mathcal{G}_{\theta}^{t:t+N}\approx\mathcal{V}_{\theta|\phi}^{t}-(1-d)\gamma^{N}\mathcal{V}_{\theta|\phi}^{t+N+1}.$ (12)
Its value gradient is then given by
$\small\nabla_{\theta}^{[q]}\mathcal{G}_{\theta}^{t:t+N}=\nabla_{\theta}\mathcal{V}_{\theta|\phi}^{t}-(1-d)\gamma^{N}\nabla_{\theta}\mathcal{V}_{\theta|\phi}^{t+N+1}$ (13)
regardless of the differentiability of the rewards. Noting that, unlike Xu et al. (2022), we specifically use action-value function $Q_{\phi}$ to compute the value to ensure $\mathcal{G}_{\theta}^{t:t+N}$ is differentiable with respect to $\theta$, which makes this derivative expression meaningful, otherwise the derivative would be zero if using $V_{\phi}$ solely with state input. Let $\nabla_{\theta}\mathcal{G}_{\theta}^{t:t+N}$ denote the first-order gradient of the accumulated reward. The average of the two gradients can be expressed as:
¯$\small\bar{\nabla}_{\theta}\mathcal{G}_{\theta}^{t:t+N}=\frac{1}{2}\left(\nabla^{[q]}_{\theta}\mathcal{G}_{\theta}^{t:t+N}+\nabla_{\theta}\mathcal{G}_{\theta}^{t:t+N}\right).$ (14)
Therefore, the difference between this gradient and the gradient (7) used in Xu et al. (2022) is that the first-order gradients in (7) are combined with the value gradients. By leveraging this combination, our method remains effective in guiding the parameter updates toward the correct direction, when the first-order gradient is biased due to the non-differentiable rewards.
We conduct a simple experiment to assess the effectiveness of incorporating the 0-step return in addressing gradient bias. We deliberately detach parts of rewards in the hovering task (see Section 4.1) to mimic non-differentiable rewards. As shown in Figure 3, combining the 0-step return with the N-step return in the objective function (9) for training significantly reduces the model parameter residuals.
We use a Gaussian policy $\pi_{\theta}(a|s)=\mathcal{N}(\mu_{\theta}(s),\sigma_{\theta}(s))$ for the actor network and apply the reparameterization trick Kingma (2013) to gradient computation. We also normalize the actions using tanh function to stabilize the training process:
$\small a_{t}=\tanh(\mu_{\theta}(s_{t})+\sigma_{\theta}(s_{t})\epsilon),\quad\epsilon\sim\mathcal{N}(0,I).$ (16)
$\small\mathcal{L}_{\phi}=\mathbb{E}_{s\in\{\tau_{i}\}}\left\|V_{\phi}(s)-\tilde{V}_{\phi}(s)\right\|^{2}.$ (17)
[…]
$\small\tilde{V}_{\phi}(s_{t})=(1-\lambda)\left(\sum_{k=1}^{N-t-1}\lambda^{k-1}G_{t}^{k}\right)+\lambda^{N-t-1}G_{t}^{N-t}$ (18)
where $G_{t}^{k}$ denotes $k-$step return from $t$:
$\small G_{t}^{k}=\left(\sum_{l=0}^{k-1}\gamma^{l}r_{t+l}\right)+(1-d)\gamma^{k}V_{\phi}(s_{t+k}).$ (19)
The state-value function is derived from the action-value function :
$\small V_{\phi}(s)=\mathbb{E}_{a\sim\pi}\left[Q_{\phi}(s,a)\right]+\kappa H\left(\pi_{\theta}(\cdot\mid s)\right),$ (20)
where we adopt an extra policy entropy term $H\left(\pi_{\theta}(\cdot\mid s)\right)$ to encourage exploration as in SAC Haarnoja et al. (2018a). $\kappa$ is an adaptive ratio whose computation follows Haarnoja et al. (2018a). To stabilize the critic training, we follow Mnih et al. (2015) to use a target critic $\phi^{-}$ to estimate the expected return (see Equation (18)).
Table 1: Observations and rewards used in benchmark quadrotor tasks
Environments Action Observation Reward Function
Hovering individual thrusts state & $\hat{p}$ $c-k_{1}\left\|p-\hat{p}\right\|-k_{2}\left\|q-\hat{q}\right\|-k_{3}\left\|v\right\|-k_{4}\left\|\omega\right\|$ (fully DIFF)
Tracking individual thrusts state & next 10 $\hat{p}_{i=1\sim 10}$ $c-k_{1}\left\|p-\hat{p}_{0}\right\|-k_{2}\left\|q-\hat{q}\right\|-k_{3}\left\|v\right\|-k_{4}\left\|\omega\right\|$ (fully DIFF)
Landing individual thrusts state & $\hat{p}$ $-k_{1}f^{+}\big{(}\left\|p_{xy}-\hat{p}_{xy}\right\|\big{)}+k_{2}f^{+}\big{(}\left\|v_{z}-\hat{v}_{z}\right\|\big{)}+k_{3}s$ (partially DIFF)
Racing individual thrusts state & next 2 $\hat{p}_{i=1,2}$ of gates $c-k_{1}\left\|p-\hat{p}_{0}\right\|-k_{2}\left\|q-\hat{q}\right\|-k_{3}\left\|v\right\|-k_{4}\left\|\omega\right\|+k_{5}s$ (partially DIFF)
  • 1
    $c$ represents a small constant used to ensure the agent remains alive. $k_{i}$ denotes constant weights for different reward contributions, with these weights being distinctly defined for each task. $s$ is a boolean variable that indicates whether the task is successfully completed. The state comprises position ($p$), orientation ($q$), linear velocity ($v$), and angular velocity ($\omega$). $f^{+}(\cdot)$ and $\hat{(\cdot)}$ are separately an increasing mapping function used to normalize the reward and target status. DIFF is abbreviation for differentiable.

### v2

@@ Abstract
Quadrotor control policies can be trained with high performance using the exact gradients of the rewards to directly optimize policy parameters via backpropagation-through-time (BPTT). However, designing a fully differentiable reward architecture is often challenging. Partially differentiable rewards will result in biased gradient propagation that degrades training performance. To overcome this limitation, we propose Amended Backpropagation-through-Time (ABPT), a novel approach that mitigates gradient bias while preserving the training efficiency of BPTT. ABPT combines 0-step and N-step returns, effectively reducing the bias by leveraging value gradients from the learned Q-value function. Additionally, it adopts entropy regularization and state initialization mechanisms to encourage exploration during training. We evaluate ABPT on four representative quadrotor flight tasks in both real world and simulation. Experimental results demonstrate that ABPT converges significantly faster and achieves higher ultimate rewards than existing learning algorithms, particularly in tasks involving partially differentiable rewards. The code will be released at Anonymous.
Keywords: Aerial Robots, Differentiable Simulation, BPTT
@@ 1 Introduction
Quadrotors have demonstrated significant potential in various real-world applications including wild rescue, dangerous high-altitude work, and delivery. Recent work [1, 2, 3] has shown end-to-end policies can be learned through imitation learning for controlling quadrotors from raw sensory data. However, the performance is largely restricted by expert’s capability. Though reinforcement learning (RL) can tackle such limitation by self-exploration of agents, it relies on RL relies on zero-order gradient (ZOG) approximations [4] which require extensive sampling or replay mechanisms, which often results in slow convergence and sub-optimal training outcomes. Compared to imitation learning and RL algorithms, recent studies [5, 6, 7, 8, 9] demonstrate that directly using first-order gradients for policy learning achieves faster convergence and superior performance, particularly for quadrotors [5, 6]. This is largely due to the good differentiable property of the quadrotor dynamics.
@@
However, using first-order gradients for training requires not only the dynamics but also the reward function to be differentiable. While this condition is manageable for simple tasks, it becomes highly challenging for complex task objectives. Reward functions in such scenarios often involve non-differentiable components, such as conditional constants or binary rewards (e.g., awarding points upon task success), which violate differentiability requirements. These non-differentiable elements disrupt the computation graph during backpropagation-through-time (BPTT), leading to biased first-order gradients. Notably, while first-order gradients remain unbiased under purely non-smooth dynamics [10, 11], they become biased when using non-differentiable rewards—a phenomenon we term Biased Gradient. This bias misguides training, causing optimization to stall in local minima and deviate from the intended direction of improvement.
Figure 1: Our trained policies were validated in real world. Two rows are separately four trials of landing (Top) and hovering (Bottom). More videos are included in supplementary materials.
@@
We propose an on-policy actor-critic approach - Amended Backpropagation-through-Time (ABPT) - to mitigate the bias gradient introduced by the non-differentiable rewards while keeping high policy learning performance in terms of training speed and converged rewards. Our approach combines 0-step returns with N-step returns [4], leveraging value gradients generated by the 0-step returns to balance first-order gradient accuracy and exploitation. Additionally, ABPT incorporates entropy to counteract the excessive aggression of first-order gradient. It also employs a replay buffer to store state experiences, initializing episodes with these states to enhance sampling efficiency. We evaluate our method on four representative quadrotor tasks, comparing it against classic policy gradient and first-order gradient methods. These tasks are designed to progressively increase the reward non-differentiability, testing the adaptability of each approach. Experimental results demonstrate that ABPT achieves the fastest convergence and highest final rewards across all baselines. This superiority is attributed to ABPT’s ability to compensate for biased gradients and enhance exploration via entropy regularization and state replay. Furthermore, ABPT exhibits robustness across varying learning rates and reward structures. Our technical contributions are summarized as follows:
@@
  • We introduce four representative quadrotor tasks as novel benchmark tests for evaluation. Our learned policies have been validated in real world as shown in Figure 1.
@@ 2 Related Work
2.1 Reinforcement Learning
Traditional reinforcement learning can be divided into two classes: model-free RL and model-based RL. Model-free RL includes value-based and policy-gradient methods. Value-based methods learn value functions to estimate long-term rewards. DQN [12] introduced neural networks for discrete actions, while DDPG [13] extended this to continuous action spaces. TD3 [14] reduced overestimation bias with multiple value networks, and SAC [15] used a maximum entropy framework for robust high-dimensional learning. Policy-gradient methods directly optimize policies using gradients. TRPO [16] stabilized updates via trust regions, and PPO [17] simplified optimization with a clipped surrogate objective.
In contrast to model-free RL which treats the environment as a black box, model-based RL [18] introduces an additional process to learn the environment’s dynamics. For example, PILCO [19] and Dyna-Q [20] leveraged learned environment models to generate simulated experiences to accelerate training. Methods like [21, 22] employ trajectory sampling to plan over learned environment models. Dreamer [23] embedded entire functions into a latent space, enabling end-to-end policy updates via backpropgation through time (BPTT). Despite their advantages, existing RL methods do not explicitly utilize the dynamics of robotics that can be precisely described by physical laws.
2.2 Differentiable Simulators
Policy learning via differentiable physics is an approach that integrates the physical simulations with differentiable dynamics to enable policy learning directly by using gradient based optimization. Making the dynamics differentiable in the simulator is the key to this approach. DiffTaichi [24] is a comprehensive differentiable physics engine that includes simulations of fluid, gas, rigid body movement, and more. In the field of robotics, Brax [25] offers differentiable versions of common RL benchmarks, built on four physics engines, including JAX and Mujoco [26]. Another line of research focuses on addressing challenges in contact-rich environments. For example, Heiden et al. [27] tackle the contact-rich discontinuity problem in quadruped robots by employing a neural network to approximate the residuals. Dojo [28] enhances contact solvers and integrates various integrators to accelerate computations while maintaining fidelity. Other research efforts focus on improving decision-making capabilities rather than control policies. VisFly [29] introduces a versatile drone simulator with fast rendering, based on Habitat-Sim [30], providing a platform for high-level applications. To enhance the efficiency, many simulators leverage GPU-accelerated frameworks like JAX [31] and PyTorch [32] for faster computations.
2.3 First-order Gradient Training
@@
With the differentiable simulators, the policy canbe trained through BPTT by using the first-order gradients. Though first-order gradients enables faster and more accurate gradient computation, accelerating policy training via BPTT [33], they suffer from gradient explosion/vanishing or instability caused by non-smooth dynamics. Many attempts try to address these issues and strengthen robustness. PODS [34] leverages both first- and second-order gradients with respect to cumulative rewards. SHAC [35] employs an actor-critic framework, truncates the learning window to avoid vanishing/exploding gradients, and smooths the gradient updates. AOBG [11] combines ZOG (policy gradient) with FOG, using an adaptive ratio based on gradient variance in the minibatch to avoid the high variance typical of pure FOG in discontinuous dynamics. AGPO [36] replaces ZOG in mixture with critic predictions, as Q-values offer lower empirical variance during policy rollouts. While both AGPO and AOBG converge to asymptotic rewards in significantly fewer timesteps, the mixture ratio requires excessive computational resources, leading to longer wall-time.
@@
Among these algorithms, SHAC [35] achieves training efficiency comparable to the naive first-order approach (BPTT). However, SHAC primarily focuses on challenges related to dynamics and overlooks the gradient bias introduced by partially differentiable rewards, which are common in real-world quadrotor tasks. These tasks often involve differentiable dynamics but rely on discrete or sparse reward signals, such as waypoint reaching events or minimum-time completions. Our approach directly tackles this limitation by introducing a 0-step return to address gradient bias from non-differentiable rewards and incorporating entropy to encourage exploration in challenging scenarios. The results highlight our method’s superior performance in quadrotor applications, achieving the fastest convergence and the highest final rewards among all the evaluated methods.
3 Preliminaries
@@
[36] named this gradient estimator as Q gradient (QG). Compared to ZOG, the performance of value approximation is extremely crucial to actor training. This is because QG directly relies on the back-propagation of the action-value function, while ZOG firstly estimates advantages among values according to current policy that enables actor training more robust to the critic.
4 First-Order Gradient Approach with Non-differentiable Rewards
@@
where $N$ represents the horizon length and $i$ denotes the $i$-th trajectory within the minibatch. To consider infinite return whiling avoiding gradient explosion, an approximated N-step return [4] has been introduced in [35]:
$\nabla_{\theta}\mathcal{J}_{\theta}=\left(\sum_{k=0}^{N-1}\gamma^{k}\frac{\partial R(s_{t+k})}{\partial\theta}\right)+\gamma^{N}\nabla_{\theta}V_{\phi}(s_{t+N}).$ (7)
Here, $V_{\phi}$ is the state-value function reparameterized by $\phi$. As shown in [35], using this approximated N-step return can introduce smooth landscape for optimization and mitigate the gradient explosion issues. However, it cannot address non-differentiable rewards as we will discuss later.
$\mathcal{J}_{\theta}=\left(\sum_{k=0}^{N-1}\gamma^{t}\Big{(}R_{dist}(s_{t+k})+R_{succ}(s_{t+k})\Big{)}\right)+\gamma^{N}V_{\phi}(s_{t+N}),$ (8)
5 The Proposed Method
@@
$\mathcal{J}_{\theta}^{N}=\underbrace{\left(\sum_{k=0}^{N-1}\gamma^{k}{R}(s_{t+k}^{i})\right)}_{\mathcal{G}^{t:t+N}_{\theta}}+(1-d)\gamma^{N}\underbrace{V_{\phi}(s_{t+N}^{i})}_{\mathcal{V}_{\theta|\phi}^{t+N+1}},\>\>\>\>\;\;\;\;\;\;\mathcal{J}_{\theta}^{0}=\underbrace{V_{\phi}(s_{t}^{i})}_{\mathcal{V}_{\theta|\phi}^{t}}$ (10)
@@
Here, $d$ is a boolean variable indicating whether the current episode has ended, and $i$ denotes the trajectory index. Given the trajectory is perform by $\pi_{\theta}$, each element is differentiable to parameter $\theta$. $\mathcal{G}_{\theta}^{t:t+N}$ represents the accumulated reward within the horizon and $\mathcal{V}_{\theta|\phi}^{t+N+1}$ is the value obtained by fixed critic. Both 0-step return and N-step return are the expected value computed from the same action-value function $Q_{\phi}$. We prove that using the objective function (9) for policy learning is equivalent to taking the average of the value gradient and the first-order gradient for backpropagation in Appendix A.
@@
We use a Gaussian policy $\pi_{\theta}(a|s)=\mathcal{N}(\mu_{\theta}(s),\sigma_{\theta}(s))$ for the actor network and apply the reparameterization trick [37] to gradient computation. We also normalize the actions using tanh function to stabilize the training process:
$a_{t}=\tanh(\mu_{\theta}(s_{t})+\sigma_{\theta}(s_{t})\epsilon),\quad\epsilon\sim\mathcal{N}(0,I).$ (11)
$\mathcal{L}_{\phi}=\mathbb{E}_{s\in\{\tau_{i}\}}\left\|V_{\phi}(s)-\tilde{V}_{\phi}(s)\right\|^{2}.$ (12)
[…]
$\tilde{V}_{\phi}(s_{t})=(1-\lambda)\left(\sum_{k=1}^{N-t-1}\lambda^{k-1}G_{t}^{k}\right)+\lambda^{N-t-1}G_{t}^{N-t}$ (13)
where $G_{t}^{k}$ denotes $k-$step return from $t$:
$G_{t}^{k}=\left(\sum_{l=0}^{k-1}\gamma^{l}r_{t+l}\right)+(1-d)\gamma^{k}V_{\phi}(s_{t+k}).$ (14)
The state-value function is derived from the action-value function :
$V_{\phi}(s)=\mathbb{E}_{a\sim\pi}\left[Q_{\phi}(s,a)\right]+\kappa H\left(\pi_{\theta}(\cdot\mid s)\right),$ (15)
where we adopt an extra policy entropy term $H\left(\pi_{\theta}(\cdot\mid s)\right)$ to encourage exploration as in SAC [15]. $\kappa$ is an adaptive ratio whose computation follows [15]. To stabilize the critic training, we follow [38] to use a target critic $\phi^{-}$ to estimate the expected return (see Equation (13)).